OpenStudy (anonymous):

Let A be a 3x4 matrix. If b=a1+a2+a3+a4. what can you conclude about the number of solutions to the linear system Ax=b?

OpenStudy (anonymous):

\[\left[\begin{matrix}x _{1} & x _{2} & x _{3} & x _{4} \\ x _{5} & x _{6} & x _{7} & x _{8} \\ x _{9} & x _{10} &x_{11} & x _{12}\end{matrix}\right] x = \left(\begin{matrix}b _{1} \\ b _{2}\\ b _{3}\\b _{4} \end{matrix}\right)\]